OpenStudy (anonymous):

Solve the system of equations by graphing. -1/3x + y = –1 y = 4 + 1/3x

OpenStudy (anonymous):

ok first rearrange the terms y=-1+1/3x y=4+1/3x

OpenStudy (anonymous):

oops no solution can this possible

OpenStudy (loser66):

no, not that, friend,

OpenStudy (anonymous):

:(

OpenStudy (loser66):

you have 2 equations , the first one is y = 1/3 x -1 (rearrange already) . so if x =0 y = -1; if y =0 , x =3. you have a graph like  and then, the second one is y = 1/3 x +4 (rearrange already), so if x =0 , y =4; if y =0 x = -12 and its graph is  @ jimmyjoew. you take out a paper, draw 2 graphs carefully , you will see they intersect at a point, and you take coordinate of that point . that's your solution